What would be a quadratic equation with 7 and -3 as its roots? How would I write the equation in the form ax^2+bx+c=0,where a, b, and c are integers?
Answer by rothauserc(4718)

we write this quadratic equation as
(x-7) * (x+3) = 0
*************************************************
note that x - 7 = 0, x = 7 and x + 3 = 0, x = -3
*************************************************
multiply (x-7) by (x+3) and we get
x^2 -4x -21 = 0
where a = 1, b = -4 and c = -21